OpenStudy (anonymous):

please help me :) Given the following system of equations x + y = 4k 2x = k where k is a constant Real Number, solve the system for (x, y) in terms of k. Do not replace the constant k with a numerical value. Give your solution as an ordered pair.

OpenStudy (anonymous):

\[\left\{\frac{k}{2},\frac{7 k}{2}\right\} \]

OpenStudy (anonymous):

what?@robtobey

OpenStudy (freckles):

You can solve it by substitution the 2x=k gives you x=k/2 replace the x in the first equation with k/2 and then solve for y

OpenStudy (anonymous):

dont understand

OpenStudy (freckles):

do you see the x=k/2 ? This means x is k/2. Do you see the x in the first equation, the equation that is x+y=4k? This means you can place that x in that equation with k/2 .

OpenStudy (anonymous):

i see that yeah and when you plug it, k/2+y=4y

OpenStudy (anonymous):

ohh okay got it

OpenStudy (freckles):

well you replace the k with y why?

OpenStudy (freckles):

like the first equation was x+y=4k not x+y=4y

OpenStudy (freckles):

you are just to replace x with k/2 and then solve for y

OpenStudy (anonymous):

okay